# A Systematic Survey on Large Language Models for Evolutionary Optimization: From Modeling to Solving

### v2

@@ Abstract
Large Language Models (LLMs), with their strong understanding and reasoning capabilities, are increasingly being explored for tackling optimization problems, especially in synergy with evolutionary computation [^1] This survey primarily focuses on evolutionary optimization, i.e., optimization based on evolutionary computation. For brevity, we use the term optimization throughout to denote this scope.. Despite rapid progress, however, the field still lacks a unified synthesis and a systematic taxonomy. This survey addresses this gap by providing a comprehensive review of recent developments and organizing them within a structured framework. We classify existing research into two main stages: LLMs for optimization modeling and LLMs for optimization solving [^2] In this survey, the terms optimization modeling and optimization solving are used as concise forms of optimization problem modeling and optimization problem solving, respectively.. The latter is further divided into three paradigms according to the role of LLMs in the optimization workflow: LLMs as stand-alone optimizers, low-level LLMs embedded within optimization algorithms, and high-level LLMs for algorithm selection and generation. For each category, we analyze representative methods, distill technical challenges, and examine their interplay with traditional approaches. We also review interdisciplinary applications spanning the natural sciences, engineering, and machine learning. By contrasting LLM-driven and conventional methods, we highlight key limitations and research gaps, and point toward future directions for developing self-evolving agentic ecosystems for optimization. An up-to-date collection of related literature is maintained at https://github.com/ishmael233/LLM4OPT.
@@ I Introduction
Optimization techniques have become indispensable for addressing complex problems across diverse domains, including engineering design [1], economic planning [2], and scientific discovery [3]. In practice, real-world challenges are typically abstracted into mathematical optimization models, which are then solved by optimization algorithms. Broadly, these algorithms can be divided into two categories: exact methods and approximate methods [4]. Exact methods, while theoretically capable of guaranteeing global optimality, often suffer from exponential growth in computational complexity, rendering them impractical for large-scale problems. In contrast, approximate methods, such as heuristics [4], can efficiently tackle large and complex problems, though without convergence guarantees. Moreover, the No Free Lunch (NFL) theorem [5] dictates that no single optimization algorithm can consistently outperform others across all problem classes. Consequently, the practical deployment of optimization algorithms typically demands substantial expertise for configuration [6], selection [7], and even customized design [8]. This heavy reliance on domain knowledge remains a critical barrier to transferring optimization technology from theory to practice.
@@
Against this backdrop, machine learning has emerged as a promising approach to alleviate the complexity of algorithm configuration and reduce dependence on expert knowledge. Among these methods, Reinforcement Learning (RL) [9, 10], with its flexible decision-making framework, has been widely applied at multiple levels of optimization algorithms [11]. At the internal structural level, RL has been employed for population initialization [12], fitness evaluation [13], operator selection [14], and parameter control [15]. At the higher strategic level, RL has enabled automated algorithm selection [16] and even algorithm generation [17]. However, a persistent limitation is that most RL-based approaches are trained on narrowly defined problem sets, which severely restricts their generalization. When confronted with new problems, retraining is often costly and time-consuming. These limitations motivate the exploration of Large Language Models (LLMs) [18], whose powerful semantic understanding, reasoning, and generalization capabilities open new opportunities for optimization research.
The rapid advancements of LLMs are fundamentally reshaping research paradigms across disciplines. Trained on vast corpora, LLMs exhibit strong capabilities in language understanding [19], complex logical reasoning [20], and open-ended knowledge generalization [21]. Their influence extends far beyond Natural Language Processing (NLP) [22], demonstrating transformative potential in high-value domains such as medicine [23] and finance [24]. For example, Huawei’s PanGu Drug Model [25] accelerates drug discovery pipelines, while BloombergGPT [26] integrates diverse financial data to provide domain-specific insights. Within optimization, the semantic and reasoning strengths of LLMs open unique opportunities: on the modeling side, they can translate natural language descriptions into mathematical optimization models [27], thereby advancing the automation and intelligence of the entire optimization workflow; on the solving side, they can function as stand-alone optimizers [28] or collaborate with existing algorithms to enhance performance; .
@@
Despite their significant potential, the use of LLMs in optimization is still at an early stage. Existing studies have already revealed important limitations. For example, LLMs as standalone optimizers often fail to surpass classical algorithms in performance [29], and although fine-tuning with domain-specific data can enhance task outcomes, constructing such datasets remains costly and expertise-intensive [27]. Against this backdrop, this survey provides a systematic analysis and guide for this emerging interdisciplinary field. The main contributions are summarized as follows:
@@
  • We propose a structured taxonomy that categorizes existing work into optimization modeling and optimization solving, with the latter further divided into three paradigms: LLMs as optimizers, low-level LLMs embedded in algorithms, and high-level LLMs for algorithm selection and generation. This taxonomy offers a unified framework for understanding and organizing the research landscape.
@@
TABLE I: Comparison of related surveys on LLMs for optimization. Our work provides the first systematic coverage of all four categories: LLMs for Modeling (LM), LLMs as Optimizers (LO), Low-level LLMs for Optimization Algorithms ($L$), and High-level LLMs for Optimization Algorithms (HL).
Ref. Venue LM LO LL HL Wu et al. [30] TEVC, 2024 Huang et al. [31] SWEVO, 2024 Chao et al. [32] Research, 2024 Yu et al. [33] arXiv, 2024 Liu et al. [34] arXiv, 2024 Ma et al. [35] TEVC, 2024 Our work arXiv, 2025
@@ II Related Surveys and Taxonomy
Several surveys and reviews have examined the use of LLMs in optimization. Table I summarizes their scope and contrasts them with our study. Wu et al. [30] and Huang et al. [31] reviewed the field from a bidirectional perspective, considering both LLM-enhanced optimization algorithms and optimization algorithm-enhanced LLMs. However, in the former direction, their discussions remained limited to scenarios where LLMs serve as solvers or for algorithm generation, without analyzing their broader relationship to the optimization process. Yu et al. [33] adopted a similar classification and proposed a general framework for LLM-assisted optimization. Chao et al. [32] focused specifically on Evolutionary Algorithms (EAs) [36, 37], highlighting component-level analogies between EAs and LLMs to suggest bidirectional knowledge transfer, but their coverage was restricted to cases where LLMs act as operators. Other related works took slightly different perspectives. For instance, Liu et al. [34] provided a broad overview of LLMs in algorithm design, categorizing their roles as optimizers, predictors, extractors, and designers. Ma et al. [35] surveyed meta-black-box optimization, which mentioned LLMs but primarily emphasized conventional reinforcement learning approaches. Overall, existing reviews have concentrated mainly on the solving stage of optimization, with limited attention to the modeling stage. Even in [34], modeling was only briefly introduced as a minor application.
In contrast to existing surveys, this paper provides a systematic review of the entire modeling-to-solving workflow of LLMs for optimization. We propose a new taxonomy that articulates the relationship between LLMs and the optimization process. Specifically, we divide the field into two main categories: LLMs for optimization modeling and LLMs for optimization solving. The latter is further subdivided into three paradigms: LLMs as optimizers, low-level LLM-assisted optimization algorithms, and high-level LLM-assisted optimization algorithms. The definitions of these categories are as follows:
@@
  • Low-level LLM-assisted Optimization Algorithms embed LLMs as intelligent components within traditional optimization algorithms to enhance specific operations. Unlike the standalone optimizer paradigm, this approach integrates LLMs tightly with established frameworks, exploiting their analytical and domain knowledge to improve performance. Typical applications include population initialization, operator design, parameter control, configuration, and fitness evaluation. By empowering traditional components to adapt more intelligently to diverse problem characteristics and search states, this paradigm improves both efficiency and solution quality at the algorithmic core.
@@ III-A Evolutionary Algorithms
Over decades of development, researchers have proposed a diverse set of classical paradigms [38, 39]. Prominent examples include Genetic Algorithms (GAs) [40], Genetic Programming (GP) [41], Differential Evolution (DE) [42], and Particle Swarm Optimization (PSO) [43]. Although these algorithms share the population-based iterative framework, they differ in search mechanisms and application domains. For instance, GAs simulate natural selection and genetic operators, making them well-suited for discrete and combinatorial problems; DE and PSO excel in continuous optimization tasks through distinctive mutation and information-sharing strategies; while GP, with its tree-structured encoding, has been widely applied to automatic programming and symbolic regression. Collectively, these paradigms established the theoretical foundation of evolutionary optimization and provided diverse strategies for tackling increasingly complex problems.
Compared to traditional exact optimization methods, EAs offer significant advantages in handling non-convex, multi-modal, and noisy problems. Their applications span a wide range of domains, including aerospace design [44] and hyperparameter tuning in deep learning [45]. Nonetheless, in line with the NFL theorem, no single EA can achieve universal superiority across all problems at low computational cost. As a result, practitioners often need to devote substantial effort to selecting, configuring, or designing suitable EA variants. To address these challenges, recent research has increasingly integrated EAs with machine learning methods [46]. Notably, reinforcement learning has been combined with evolutionary frameworks, giving rise to approaches such as MOEA/D-DQN [47] and SYMBOL [17]. With the emergence of LLMs, cutting-edge studies are further exploring how LLMs can empower EAs [30], extending from problem formulation [27] to automated algorithm design [34].
@@ III-B Large Language Models
The foundation of LLMs lies in the Transformer architecture [48], which captures global input-output dependencies through self-attention. Building on this architecture, researchers have developed three primary paradigms: encoder-decoder, encoder-only, and decoder-only models [49]. Encoder-decoder models, exemplified by BART [50] and T5 [51], follow the original Transformer design and excel at sequence-to-sequence tasks. Encoder-only models, such as BERT [52], focus on contextual representation learning for semantic understanding. Decoder-only models, including ChatGPT [53] and DeepSeek [54], leverage autoregressive generation for open-domain tasks and currently dominate LLM-for-Optimization research due to their strong generative and reasoning capabilities. Relevant applications mainly employ two pathways: prompt engineering [55] and fine-tuning [56].
Prompt engineering steers LLM outputs by carefully designed instructions without modifying model parameters. Early forms include zero-shot and few-shot prompting [57], which exploit pre-trained knowledge directly or via limited in-context examples. More advanced approaches incorporate structured reasoning, such as Chain-of-Thought (CoT) [20], which enhances performance on complex reasoning tasks by eliciting step-by-step logical inference. CoT has since expanded into richer paradigms like Tree-of-Thought (ToT) [58], Graph-of-Thought (GoT) [59], and domain-specific variants such as Chain-of-Code (CoC) [60] and Chain-of-Knowledge (CoK) [61].
Fine-tuning, in contrast, updates model parameters on task-specific data, enabling deeper adaptation at the cost of labeled data and computational resources. Two strategies are particularly relevant. Instruction fine-tuning [62] reformulates diverse tasks into instruction–input–output triplets, unifying them under an instruction-following paradigm. This significantly improves zero-shot generalization, as demonstrated by Flan-T5 [63]. Alignment fine-tuning ensures model behavior aligns with human intentions and values. Reinforcement Learning with Human Feedback (RLHF) [64] achieves this by optimizing a reward model built from human preferences, while Direct Preference Optimization (DPO) [65] simplifies the process by directly optimizing for preferred outputs. Together, these techniques are central to building safe, reliable, and useful LLMs.
@@ IV LLMs for Optimization Modeling
Optimization modeling is a crucial step in the optimization workflow but traditionally relies heavily on expert knowledge and experience. Recently, LLMs have opened new opportunities for automated optimization modeling. In this section, we introduce two core paradigms: prompt-based methods and learning-based methods. Prompt-based methods (Section IV-A) guide LLMs through carefully designed instructions, often implemented via two-stage workflows, multi-agent collaboration, or interactive prompting. Learning-based methods (Section IV-B), in contrast, fine-tune LLMs to directly generate mathematical formulations, typically requiring large-scale synthetic data for training but offering greater reliability in complex tasks. Finally, we discuss open challenges in this field (Section IV-C). Representative studies are summarized in Table II, while the main paradigms are illustrated in Fig. 3.
@@
Method Venue Type Technical Summary Prompt-based Methods Ner4OPT [66] CPAIOR, 2023 Two-stage Fine-tune models for named entity recognition by incorporating traditional NLP methods. AOMG [67] GECCO, 2023 Direct Directly leverage LLMs to generate mathematical models. HG 2.0[68] arXiv, 2023 Two-stage Embed LLMs within the two-stage framework. AMGPT [69] arXiv, 2023 Two-stage Employ fine-tuned models for constraint classification. CoE [70] ICLR, 2023 Multi-agent Construct dynamic reasoning chains using 11 expert agents. OptiMUS [71] ICML, 2023 Multi-agent Develop a conductor agent to coordinate modeling, programming, and evaluation processes. MAMS [72] INFOR, 2024 Multi-agent Develop inter-agent cross-validation to replace solver-dependent verification. EC [73] SGC, 2024 Two-stage Apply the two-stage framework to energy management systems. CAFA [74] NeurIPS, 2024 Two-stage Enhance modeling performance through code-based formalization. MAMO [75] NAACL, 2024 Two-stage Develop the MAMO benchmark with ordinary differential equation extensions. NL2OR [76] arXiv, 2024 Two-stage Predefine abstract structural constraints to regulate the LLM outputs. TRIP-PAL [77] arXiv, 2024 Two-stage Apply the two-stage framework to travel planning. OptLLM [78] arXiv, 2024 Interactive Support both single-input and interactive-input modes. OptiMUS-0.3 [79] arXiv, 2024 Multi-agent Introduce self-corrective prompts and structure-aware modeling based on OptiMUS. MeetMate [80] TiiS, 2024 Interactive Develop an interactive system for user input processing with five selectable task options. VL [81] CUI, 2025 Interactive Convert user priorities into optimization constraints during dialog interactions. LLM-MCTS [82] arXiv, 2025 Two-stage Conduct hierarchical Monte Carlo tree search over the hypothesis space. EquivaMap [83] arXiv, 2025 Two-stage Generate variable mapping functions via LLMs with lightweight verification. MAP [84] arXiv, 2025 Multi-agent Deploy multiple independent reviewers to evaluate modeling results. ORMind [85] arXiv, 2025 Multi-agent Replace the conductor agent with structured, predictable workflows. LEAN-LLM-OPT [86] ssrn, 2025 Multi-agent Integrate RAG to enhance problem classification, and construct instances with query operations. Learning-based Methods LM4OPT [87] INFOR, 2024 Fine-tuning Progressively fine-tune models on the NL4OPT dataset. ORLM [27] OR, 2024 Data synthesis Synthesize data through expansion and augmentation and fine-tune open-source models. ReSocratic [88] ICLR, 2024 Data synthesis Propose inverse data synthesis methodology and construct the OPTIBENCH benchmark. LLMOPT [89] ICLR, 2024 Fine-tuning Introduce model alignment and self-correction mechanisms to mitigate hallucination phenomena. BPP-Search [90] arXiv, 2024 Data synthesis Solve the problem of missing detailed in data synthesis. OptMATH [91] arXiv, 2025 Data synthesis Develop a scalable bidirectional data synthesis approach. LLMBO [92] arXiv, 2025 Data synthesis Propose a method for fine-tuning cost-effective LLMs to tackle specific business challenges. SIRL [93] arXiv, 2025 Fine-tuning Integrate external optimization solvers as verifiable reward validators for reinforcement learning. Step-Opt [94] arXiv, 2025 Data synthesis Increase problem complexity through an iterative problem generation approach. DPLM [95] arXiv, 2025 Data synthesis Combine the diversity of forward generation and the reliability of inverse generation. OptiTrust [96] arXiv, 2025 Data synthesis Develop a verifiable synthetic data generation pipeline.

### v3

@@ Abstract
Large Language Models (LLMs), with their strong understanding and reasoning capabilities, are increasingly being explored for tackling optimization problems, especially in synergy with evolutionary computation [^1] This survey primarily focuses on evolutionary optimization, i.e., optimization based on evolutionary computation. For brevity, we use the term optimization throughout to denote this scope.. While several recent surveys have explored aspects of LLMs for optimization, there remains a need for an integrative perspective that connects problem modeling with solving workflows. This survey addresses this gap by providing a comprehensive review of recent developments and organizing them within a structured framework. We classify existing research into two main stages: LLMs for optimization modeling and LLMs for optimization solving [^2] In this survey, the terms optimization modeling and optimization solving are used as concise forms of optimization problem modeling and optimization problem solving, respectively.. The latter is further divided into three paradigms according to the role of LLMs in the optimization workflow: LLMs as stand-alone optimizers, low-level LLMs embedded within optimization algorithms, and high-level LLMs for algorithm selection and generation. For each category, we analyze representative methods, distill technical challenges, and examine their interplay with traditional approaches. We also review interdisciplinary applications spanning the natural sciences, engineering, and machine learning. By contrasting LLM-driven and conventional methods, we highlight key limitations and research gaps, and point toward future directions for developing self-evolving agentic ecosystems for optimization. An up-to-date collection of related literature is maintained at https://github.com/ishmael233/LLM4OPT.
@@ I Introduction
Optimization techniques have become indispensable for addressing complex problems across diverse domains [chen2022evolutionary, cambria2014computational], including engineering design [chai2013evolutionary], economic planning [mahor2009economic], and scientific discovery [terayama2021black]. In practice, real-world challenges are typically abstracted into mathematical optimization models, which are then solved by optimization algorithms. Broadly, these algorithms can be divided into two categories: exact methods and approximate methods [festa2014brief]. Exact methods, while theoretically capable of guaranteeing global optimality, often suffer from exponential growth in computational complexity, rendering them impractical for large-scale problems. In contrast, approximate methods, such as heuristics [festa2014brief], can efficiently tackle large and complex problems, though without convergence guarantees. Moreover, the No Free Lunch (NFL) theorem [wolpert2002no] dictates that no single optimization algorithm can consistently outperform others across all problem classes. Consequently, the practical deployment of optimization algorithms typically demands substantial expertise for configuration [hutter2009paramils], selection [kerschke2019automated], and even customized design [muller1981heuristics]. This heavy reliance on domain knowledge remains a critical barrier to transferring optimization technology from theory to practice.
@@
Against this backdrop, machine learning [zhang2011evolutionary] has emerged as a promising approach to alleviate the complexity of algorithm configuration and reduce dependence on expert knowledge. Among these methods, Reinforcement Learning (RL) [kaelbling1996reinforcement, li2017deep], with its flexible decision-making framework, has been widely applied at multiple levels of optimization algorithms [mazyavkina2021reinforcement]. At the internal structural level, RL has been employed for population initialization [mundhenk2021symbolic], fitness evaluation [wang2022surrogate], operator selection [pettinger2002controlling], and parameter control [eiben2006reinforcement]. At the higher strategic level, RL has enabled automated algorithm selection [lagoudakis2000algorithm] and even algorithm generation [chen2024symbol]. However, a persistent limitation is that most RL-based approaches are trained on narrowly defined problem sets, which severely restricts their generalization. When confronted with new problems, retraining can be costly and time-consuming.
These limitations motivate exploration of Large Language Models (LLMs) [zhao2023survey, yan2024generative], which bring a distinct and complementary set of capabilities. Trained on massive, heterogeneous corpora, LLMs demonstrate strong semantic understanding, compositional abstraction, and in‑context learning that enable zero or few‑shot transfer across heterogeneous tasks [patil2024review, wei2022chain, feng2024don]. Such strengths make LLMs particularly useful when objectives, constraints, or design criteria are naturally expressed in language or structured symbols, and when rapid, readable candidate generation, explanation, or cross‑domain priors are valuable (e.g., drug discovery [mi2022pangu], finance [wu2023bloomberggpt]).
Within optimization, these capabilities open opportunities on both the modeling and solving sides: on the modeling side, LLMs can assist in translating informal specifications into formal mathematical formulations [huang2025orlm]; on the solving side, they can be used as stand‑alone iterative solvers in language‑structured problems [yang2023large] or as components that complement algorithmic search [liu2024evolution]. Despite their significant potential, the use of LLMs in optimization is still at an early stage. Existing studies have already revealed important limitations. For example, LLMs as stand-alone optimizers often fail to surpass classical algorithms in performance [huang2024exploring], and although fine-tuning with domain-specific data can enhance task outcomes, constructing such datasets remains costly and expertise-intensive [huang2025orlm].
@@
Against this backdrop, this survey provides a systematic analysis and guide for this emerging interdisciplinary field. The main contributions are summarized as follows:
@@
  • We propose a structured taxonomy that categorizes existing work into optimization modeling and optimization solving, with the latter further divided into three paradigms: LLMs as optimizers, low-level LLMs embedded in algorithms, and high-level LLMs for algorithm selection and generation. This taxonomy provides an organizing lens to help compare methods and understand the research landscape.
@@ II Related Surveys and Taxonomy
Several surveys and reviews have examined the use of LLMs in optimization. Table I summarizes their scope and contrasts them with our study. Wu et al. [wu2025evolutionary] and Huang et al. [huang2024large] reviewed the field from a bidirectional perspective, considering both LLM-enhanced optimization algorithms and optimization algorithm-enhanced LLMs. However, in the former direction, their discussions remained limited to scenarios where LLMs serve as solvers or for algorithm generation, without analyzing their broader relationship to the optimization process. Yu et al. [yu2024deep] adopted a similar classification and proposed a general framework for LLM-assisted optimization. Chao et al. [chao2024large] focused specifically on Evolutionary Algorithms (EAs) [bartz2014evolutionary, fonseca1995overview], highlighting component-level analogies between EAs and LLMs to suggest bidirectional knowledge transfer, but their coverage was restricted to cases where LLMs act as operators. Other related works took slightly different perspectives. For instance, Liu et al. [liu2024systematic] provided a broad overview of LLMs in algorithm design, categorizing their roles as optimizers, predictors, extractors, and designers. Ma et al. [ma2025toward] surveyed meta-black-box optimization, which mentioned LLMs but primarily emphasized conventional reinforcement learning approaches. Overall, existing reviews have concentrated mainly on the solving stage of optimization, with limited attention to the modeling stage. Even in [liu2024systematic], modeling was only briefly introduced as a minor application.
TABLE I: Comparison of related surveys on LLMs for optimization. This work provides a systematic coverage emphasizing all four categories: LLMs for Modeling (LM), LLMs as Optimizers (LO), Low-level LLMs for Optimization Algorithms ($L$), and High-level LLMs for Optimization Algorithms (HL).
Ref. Venue LM LO LL HL Wu et al. [wu2025evolutionary] TEVC, 2024 Huang et al. [huang2024large] SWEVO, 2024 Chao et al. [chao2024large] Research, 2024 Yu et al. [yu2024deep] arXiv, 2024 Liu et al. [liu2024systematic] arXiv, 2024 Ma et al. [ma2025toward] TEVC, 2024 Our work arXiv, 2025
Building on existing surveys, this paper focuses on the modeling-to-solving workflow and attempts a more integrative view that links modeling and solving perspectives. We propose a new taxonomy that articulates the relationship between LLMs and the optimization process. Specifically, we divide the field into two main categories: LLMs for optimization modeling and LLMs for optimization solving. The latter is further subdivided into three paradigms: LLMs as optimizers, low-level LLM-assisted optimization algorithms, and high-level LLM-assisted optimization algorithms. The definitions of these categories are as follows:
@@
  • Low-level LLM-assisted Optimization Algorithms embed LLMs as intelligent components within traditional optimization algorithms to enhance specific operations. Unlike the stand-alone optimizer paradigm, this approach integrates LLMs tightly with established frameworks, exploiting their analytical and domain knowledge to improve performance. Typical applications include population initialization, operator design, parameter control, configuration, and fitness evaluation. By empowering traditional components to adapt more intelligently to diverse problem characteristics and search states, this paradigm improves both efficiency and solution quality at the algorithmic core.
@@ III-A Evolutionary Algorithms
Over decades of development, researchers have proposed a diverse set of classical paradigms [cheng2014competitive, tian2017platemo]. Prominent examples include Genetic Algorithms (GAs) [holland1992genetic], Genetic Programming (GP) [koza1994genetic], Differential Evolution (DE) [das2010differential], and Particle Swarm Optimization (PSO) [kennedy1995particle]. Although these algorithms share the population-based iterative framework, they differ in search mechanisms and application domains. For instance, GAs simulate natural selection and genetic operators, making them well-suited for discrete and combinatorial problems; DE and PSO excel in continuous optimization tasks through distinctive mutation and information-sharing strategies; while GP, with its tree-structured encoding, has been widely applied to automatic programming and symbolic regression. Collectively, these paradigms established the theoretical foundation of evolutionary optimization and provided diverse strategies for tackling increasingly complex problems.
Compared to traditional exact optimization methods, EAs offer significant advantages in handling non-convex, multi-modal, and noisy problems. Their applications span a wide range of domains, including aerospace design [arias2012multiobjective] and hyperparameter tuning in deep learning [elsken2019neural]. Nonetheless, in line with the NFL theorem, no single EA can achieve universal superiority across all problems at low computational cost. As a result, practitioners often need to devote substantial effort to selecting, configuring, or designing suitable EA variants. To address these challenges, recent research has increasingly integrated EAs with machine learning methods [talbi2021machine]. Notably, reinforcement learning has been combined with evolutionary frameworks, giving rise to approaches such as MOEA/D-DQN [tian2022deep] and SYMBOL [chen2024symbol]. With the emergence of LLMs, cutting-edge studies are further exploring how LLMs can empower EAs [wu2025evolutionary], extending from problem formulation [huang2025orlm] to automated algorithm design [liu2024systematic].
@@ III-B Large Language Models
The foundation of LLMs lies in the Transformer architecture [vaswani2017attention], which captures global input-output dependencies through self-attention. Building on this architecture, researchers have developed three primary paradigms: encoder-decoder, encoder-only, and decoder-only models [cai2021compare]. Encoder-decoder models, exemplified by BART [lewis2019bart] and T5 [raffel2020exploring], follow the original Transformer design and excel at sequence-to-sequence tasks. Encoder-only models, such as BERT [devlin2019bert], focus on contextual representation learning for semantic understanding. Decoder-only models, including ChatGPT [floridi2020gpt] and DeepSeek [guo2024deepseek], leverage autoregressive generation for open-domain tasks and currently dominate LLM-for-Optimization research due to their strong generative and reasoning capabilities. Relevant applications mainly employ two pathways: prompt engineering [white2023prompt] and fine-tuning [han2024parameter].
Prompt engineering steers LLM outputs by carefully designed instructions without modifying model parameters. Early forms include zero-shot and few-shot prompting [kojima2022large], which exploit pre-trained knowledge directly or via limited in-context examples. More advanced approaches incorporate structured reasoning, such as Chain-of-Thought (CoT) [wei2022chain], which enhances performance on complex reasoning tasks by eliciting step-by-step logical inference. CoT has since expanded into richer paradigms like Tree-of-Thought (ToT) [yao2023tree], Graph-of-Thought (GoT) [besta2024graph], and domain-specific variants such as Chain-of-Code (CoC) [li2023chain] and Chain-of-Knowledge (CoK) [li2023chain2].
Fine-tuning, in contrast, updates model parameters on task-specific data, enabling deeper adaptation at the cost of labeled data and computational resources. Two strategies are particularly relevant. Instruction fine-tuning [liu2023visual] reformulates diverse tasks into instruction–input–output triplets, unifying them under an instruction-following paradigm. This significantly improves zero-shot generalization, as demonstrated by Flan-T5 [chung2024scaling]. Alignment fine-tuning ensures model behavior aligns with human intentions and values. Reinforcement Learning with Human Feedback (RLHF) [ouyang2022training] achieves this by optimizing a reward model built from human preferences, while Direct Preference Optimization (DPO) [rafailov2023direct] simplifies the process by directly optimizing for preferred outputs. Together, these techniques are central to building safe, reliable, and useful LLMs.
@@ IV LLMs for Optimization Modeling
Optimization modeling is a crucial step in the optimization workflow but traditionally relies heavily on expert knowledge and experience. Recently, LLMs have opened new opportunities for automated optimization modeling. In this section, we introduce two core paradigms: prompt-based methods and learning-based methods. Prompt-based methods (Section IV-A) guide LLMs through carefully designed instructions, often implemented via two-stage workflows, multi-agent collaboration, or interactive prompting. Learning-based methods (Section IV-B), in contrast, fine-tune LLMs to directly generate mathematical formulations, typically requiring large-scale synthetic data for training but offering greater reliability in complex tasks. Finally, we discuss open challenges in this field (Section IV-C). Representative studies are summarized in Table I in Supplementary Document, while the main paradigms are illustrated in Fig. 3.